A company makes snow globes each snow globe is in the shape of a sphere and the diameter of the sphere is 8 centimeters one cubic centimeters of water has a mass of about 1gram if the snow globe is filled with water what is the total mass of the water in the snow globe use 3.14 and round to the nearest whole gram
Answer by Boreal(15235)

V=(4/3)*pi *r^3, where r is half the diameter or 4 cm, and 4^3=64 cm^3
V=(4/3)*64*3.14=267.95 cm^3 rounded to 268 cm^3; since 1 cm^3 water has 1 gm weight, the total water is 268 gm.